OpenStudy (anonymous):

can anybody solve this problem for me?? 4(24/8)+(11x2)=?

OpenStudy (nowhereman):

At which point is your problem?

OpenStudy (anonymous):

the entire equation, i need the answer to prove a point,but i really dont understand algebra

OpenStudy (nowhereman):

Just evaluate from inner to outer terms. You should be able to do basic operations with numbers.

OpenStudy (anonymous):

4 * 24 11 * 2 -- + 8 4 goes into 8 2 times 1 * 24 11 * 2 -- + 2 You do division and multiplication before addition 24/2 = 12 11 * 2 = 22 12 + 22 = 34